OpenStudy (anonymous):

4 + 2 [8 + (3 - 1)^2] my answer is 72

OpenStudy (raden):

unlucky :) 3-1=2 2^2=4 8+4=12 2 * 12 = 24 4 + 24 = 28

OpenStudy (anonymous):

how did we get from all your hard questions to arithmetic?

OpenStudy (anonymous):

3 - 1 = 2 2^2

OpenStudy (anonymous):

@satellite73 switched courses

OpenStudy (anonymous):

2^2= 4

OpenStudy (anonymous):

8 + 4? I thought?

OpenStudy (phi):

yes

OpenStudy (anonymous):

8 + 4 = 12

OpenStudy (anonymous):

12 + 4 + 2 12 + 6 = 18

OpenStudy (phi):

oops

OpenStudy (anonymous):

at which part?

OpenStudy (phi):

4 + 2 [12]

OpenStudy (anonymous):

ohhhhhh brackets

OpenStudy (phi):

you have 4 plus 2 times 12 multiplication first

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i appreciate your help so much

OpenStudy (phi):

after you get to 4 + 2[12] it is not the brackets... it is that you are doing multiplication

OpenStudy (phi):

in other words 2*(12) is the same as 2*12 (we can drop the parens)

OpenStudy (phi):

if you want to see videos http://www.khanacademy.org/math/arithmetic/multiplication-division/order_of_operations/v/introduction-to-order-of-operations and the others in that section might make all this very clear. Btw, it makes no sense to not learn this, because if you don't the rest of math will be a total mystery... you need to know this to do more complicated stuff.

OpenStudy (anonymous):

it was one simple mistake...jesus christ...thanks